Approximately 5 out of every 10 students at Kendra's college live on campus. The college has 5,000 students. How many of them live on campus?
Answer by jojo14344(1513)

Well, as it stated on the problem that 5 for every 10 live in campus, and that is 50% (5/10=1/2=.50=50%). So if there are 5,000 college students, 50% of that=
5,000*.50= 2,500 of them live in campus.
Thank you,
Jojo